OpenStudy (anonymous):

Garrett throws dart at a circular dart board. the dart board has a radius of 16 inches, and the bull's eye in the center of the dart has a radius of 6 inches. What is the probability that a dart thrown at random within the dartboard will hit the bull's eye?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

i got 14.1% is that right?

OpenStudy (anonymous):

I get around 16.3% but that isn't one of my answer choices

OpenStudy (kropot72):

The probability is found by dividing the area of the bull's eye by the total area of the dart board: \[\large P(bull's\ eye)=\frac{\pi\ 6^{2}}{\pi\ 16^{2}}=\frac{36}{256}=0.1406\] Therefore your answer of 14.1% is close.